# Michael Callen Miguel Fajardo-Steinhäuser Michael G. Findley

### v2

@@ 4 Experts’ beliefs and methodological safeguards
We measure the expert views by surveying 55 international analysts, practitioners and policymakers working on global development issues (including many engaged specifically on hunger in Afghanistan) and 36 academics, many with expertise related to digital payment systems. [^23] The policy organizations represented include the USAID, the UN World Food Program, UNICEF, the World Bank, GiveDirectly, Brookings and BRAC, and the academic institutions included Brown, Duke, Georgetown, the London School of Economics, Princeton, Stanford, the University of Chicago, UC Berkeley, and UC San Diego. The survey provided a precise description of the intervention and asked respondents to predict the corresponding treatment impacts for several key outcomes we report here; see Table A7 for the full text of the survey prompt.
Table A8 compares the actual values of the intervention’s components (column 1) to the respondents’ predicted values (column 2), and shows the $p$-value of a test of equality (column 3). Across all four components, differences between expert predictions and the values from our data are statistically significant. These respondents predicted, on average, that only 43% of the women in our study would be able to use digital payments (SD = 25.67pp). In practice, as described above, 99.75% of women purchased goods at least once. They also predicted, on average, that local authorities would attempt to tax roughly 40% of beneficiaries (SD = 27.28). In practice, less than 2% of the treatment group reported any diversion attempts, with no statistically significant difference in the control group. Moreover, respondents had little agreement in their views, reflected in the large variance across predictions. A combination of skepticism and uncertainty about potential efficacy might partly explain an unwillingness to adopt innovations, especially when stakes are high. Experts did a better job of predicting cost-efficiency, guessing on average that delivery costs would be 10.65 cents on the dollar, while the actual delivery costs were 6.7 cents on the dollar, indicating a more common belief that digital aid delivery is relatively cheap.
@@ 5 Discussion
We expect that whether the findings obtained from this context and population generalize to other settings will depend most critically on three key enabling conditions. First, vulnerable households must have access to phones. Second, there must be enough merchants who accept digital payments to allow convenience and competitive shopping. And third, markets must have enough goods to meet demand. Transitioning from in-kind to cash and digital support can increase competition for goods and so may increase market prices and thereby displace hunger to other vulnerable groups. Studies of famine highlight that in extreme cases of sudden and widespread deprivation – as in Gaza at the time of writing (Chotiner, 2024) – increasing food supply is critical, but that response to most famines historically required complementary strategies to redistribute resources (Dreze and Sen, 1990; Sen, 1999).
@@
Based largely on this research, humanitarian cash-based programming is growing and digital delivery is increasingly popular (Urquhart et al., 2023), but research during humanitarian crises with populations facing extreme cultural and political constraints is still nascent. In Online Appendix 1, we review the literature and find 23 RCTs of cash for humanitarian applications, 10 such RCTs in contexts categorized as “not free” by Freedom House, three of those 10 having some digital component, and none, that we can find, targeting populations that the government actively oppresses. This knowledge gap is consequential as these contexts and groups are increasingly affected by acute hunger (Foreign, Commonwealth and Development Office, 2023). As our expert survey demonstrates, our results are not readily inferred from prior studies: illiterate Afghan women could have been unable to utilize digital payments, local merchants could have refused to accept digital transactions, and the Taliban authorities could have interfered. Our study thus provides an “existence result" on the feasibility of digital aid in humanitarian response, with insight into the organizational and technological factors that contributed to success.
@@
The exigencies of this context and the fact that we work with a hard-to-reach group, however, necessitated three design choices. First, our experimental data cover only two months before the early group is treated. This time frame, however, is consistent with our learning objective: can digital payments cost-effectively reduce acute hunger. As such, we pre-specified this as the relevant time frame, and relatedly that we did not expect changes in outcomes like income or employment. Such impacts were unlikely – and did not realize – given the large humanitarian needs, prohibitions on women’s employment, and relatively modest transfer size. Second, working with our sample required approval of the de facto authorities. Our partners could not justify seeking Taliban permission to onboard women to the digital payments app without eventually providing humanitarian assistance. We thus opted for a short, staggered program where every participant received some assistance during the lean season. While the control group knew they would receive transfers, they were unaware when these would start and we find no evidence of borrowing or other anticipatory changes in behavior. Our late group, thus, provides a valid counterfactual. Third, because we could not visit our participants, and to avoid survey fatigue over the phone, we focused on a small set of questions and could not take measurements that would yield deeper insights (such as anthropometrics). Thus, other relevant topics (e.g., within household dynamics, validated mental health measures, and impacts on other household members) were not covered.
@@
Nonetheless, this exercise provides proof-of-concept that digital aid represents a potential cost-effective complement to existing modalities. We see five important directions for future research to help humanitarians select strategies to address hunger. First, relative to physical goods and cash, digital payments enable humanitarians to more easily vary when aid is delivered, its frequency, and the amount. Research can help identify how to optimally take advantage of this added flexibility. Second, more evidence is needed on the costs and benefits of digital relative to cash delivery. Some prior work addresses this question (Aker et al., 2016), but we think more of this work in fragile contexts controlled by hostile actors and where the vulnerable are hard-to-reach will be relevant given the rise in hunger among such groups. This is best pursued in experiments which randomize both digital delivery and cash delivery against a control group. Third, humanitarians need further evidence on how key enabling conditions – such as mobile phone penetration and the availability of merchants who accept digital payments – affect these relative costs and benefits, and how best to organize future hybrid partnerships based on differences in enabling conditions. Fourth, additional focus is needed on approaches to targeting and onboarding beneficiaries in hostile environments that minimize risks to participants (Aiken et al., 2022; Jeong and Trako, 2022). In particular, the one point of physical contact in our study between humanitarians and beneficiaries was during onboarding. Further innovations might eliminate the need for contact between humanitarians and those in need altogether. And fifth, implementing digital payments at scale may affect diversion strategies, market prices, and the broader use of digital payments technologies for other financial transactions. It may also spur the transformation of humanitarian organizations into leaner organizational models to best complement the capacities of private and community organizations. Future work with large-scale experiments should explore such changes.
Encouragingly, aid agencies are increasing their use of digital payments. The WFP is using insights from our study to scale digital payments to over 100,000 households in Afghanistan. The existence of key enablers for the success of a program like this in other fragile settings (e.g., the presence of CDCs, mobile phone availability, and digital payments networks), as shown in Table A9, suggests that this approach could increase food security not only in Afghanistan but also in other fragile states with high levels of hunger.
@@
Control Control Treatment Standard Naive Adjusted Mean SD Effect Error p-value p-value N (1) (2) (3) (4) (5) (6) (7) Panel A. Primary Outcomes Days skipping meals (past week) 2.569 1.717 –0.76 0.051 0 0.0002 4761 Children skipping meals (=1) 0.873 0.333 –0.117 0.012 0 0.0002 4750 Regularly eat twice a day 0.501 0.5 0.093 0.015 0 0.0002 4736 Total bread and tea meals (past week) 13.639 3.549 –1.608 0.121 0 0.0002 4763 Food Security - KLK Index 0 1 0.501 0.032 0 4763 Better economic situation 0.048 0.213 0.335 0.011 0 0.0002 4762 Satisfied with fin. situation 0.133 0.34 0.263 0.012 0 0.0002 4755 Happy 0.154 0.361 0.28 0.014 0 0.0002 4732 Life satisfaction 3.179 1.707 1.963 0.068 0 0.0002 4763 Economic/Wellbeing - KLK Index 0 1 1.498 0.042 0 4763 Panel B. Secondary Outcomes Days eating rice (past week) 0.698 1.058 0.595 0.035 0 0.001 4763 Days eating beans (past week) 0.518 0.836 0.493 0.029 0 0.001 4763 Days eating vegetables (past week) 1.394 1.346 –0.003 0.041 0.936 0.454 4763 Days eating chicken (past week) 0.021 0.147 0.012 0.006 0.035 0.037 4763 Days eating dairy (past week) 0.074 0.383 0.047 0.013 0 0.001 4763 Able to buy medicine 0.051 0.221 0.035 0.01 0.001 0.002 3582 Involved in fin. decisions 0.666 0.472 0.017 0.015 0.271 0.184 4757 Total household income (past month) 876.683 1581.531 129.119 91.029 0.156 0.117 4763 Household’s head employed (past month) 0.192 0.394 –0.008 0.014 0.59 0.306 4741
@@
(1) (2) (3) (4) (5) Panel A. Food Security Days Skipping Children Skipping Regularly Eat Total Bread KLK Index Meals Meals (=1) Twice a Day (=1) & Tea Meals $\beta_{1}$: Treated $\times$ Round 2 –0.580*** –0.096*** –0.006 –0.949*** 0.309*** (0.094) (0.021) (0.028) (0.219) (0.056) $\beta_{2}$: Round 2 –0.194*** 0.025* 0.072*** –0.321** 0.096** (0.071) (0.013) (0.020) (0.151) (0.038) $\beta_{3}$: Treated –0.470*** –0.069*** 0.096*** –1.133*** 0.346*** (0.072) (0.015) (0.020) (0.163) (0.042) Control Mean 2.569 0.873 0.501 13.639 0.000 Observations 4,761 4,750 4,736 4,763 4,763 ($\beta_{1}$ + $\beta_{2}$ + $\beta_{3}$) / $\beta_{3}$ 2.650 2.040 1.700 2.120 2.170 p-value $\beta_{1}$ + $\beta_{3}$ = 0 0 0 0 0 0 Panel B. Wellbeing Better Economic Satisfied with Happy Life Satisfaction KLK Index Situation Financial Sit. $\beta_{1}$: Treated $\times$ Round 2 0.219*** 0.200*** 0.102*** 0.842*** 0.837*** (0.021) (0.024) (0.026) (0.127) (0.078) $\beta_{2}$: Round 2 –0.013 –0.036*** 0.030* –0.223*** –0.074* (0.009) (0.014) (0.017) (0.075) (0.040) $\beta_{3}$: Treated 0.225*** 0.163*** 0.229*** 1.542*** 1.079*** (0.015) (0.017) (0.019) (0.093) (0.055) Control Mean 0.048 0.133 0.154 3.179 0.000 Observations 4,762 4,755 4,732 4,763 4,763 ($\beta_{1}$ + $\beta_{2}$ + $\beta_{3}$) / $\beta_{3}$ 1.910 2.010 1.580 1.400 1.710 p-value $\beta_{1}$ + $\beta_{3}$ = 0 0 0 0 0 0
$Notes$: This table reports estimated impacts of treatment separately for the first and second survey round. Households were surveyed once per month for two months. Each of these months constitutes a survey round. All specifications control for stratum fixed effects and the baseline values of the dependent variables, if available. Standard errors are clustered at the individual level. The outcome variables follow the primary outcomes shown in Table 1.
@@
Gov. Off. Comm. Leader Gov. Off. Comm. Leader KLK Yes to Others Others You You Index Any Question (1) (2) (3) (4) (5) (6) Panel A. Baseline Treated 0.001 0.002 0.002 0.002 0.074* 0.003 (0.001) (0.002) (0.001) (0.003) (0.042) (0.003) Observations 4,558 4,596 4,631 4,626 4,648 4,509 Control Mean 0.003 0.002 0.001 0.005 0.000 0.009 Panel B. Long-Run $\beta_{1}$: Treated $\times$ Round 2 –0.000 0.006* 0.002 –0.001 0.092 0.004 (0.003) (0.003) (0.003) (0.005) (0.084) (0.006) $\beta_{2}$: Round 2 –0.002 –0.003 –0.003* –0.008** –0.138*** –0.010** (0.002) (0.002) (0.001) (0.003) (0.045) (0.004) $\beta_{3}$: Treated 0.001 –0.007 –0.001 0.003 –0.064 –0.003 (0.005) (0.005) (0.005) (0.009) (0.139) (0.011) p-value $\beta_{1}$ + $\beta_{3}$ = 0 0.699 0.698 0.721 0.585 0.661 0.857 ($\beta_{1}$ + $\beta_{2}$ + $\beta_{3}$) / $\beta_{3}$ –0.850 0.500 1.810 –1.670 1.700 2.600 Observations 4,558 4,596 4,631 4,626 4,648 4,509 Panel C. List Experiment All Sample Late Sample Early Sample Treated List 0.011 0.019 0.008 (0.023) (0.032) (0.032) Observations 2,358 1,171 1,187 Mean Items Mentioned 0.746 0.691 0.800

### v4

@@ 4 Experts’ beliefs and methodological safeguards
We measure the expert views by surveying 55 international analysts, practitioners and policymakers working on global development issues (including many engaged specifically on hunger in Afghanistan) and 36 academics, many with expertise related to digital payment systems. [^26] The policy organizations represented include the USAID, the UN World Food Program, UNICEF, the World Bank, GiveDirectly, Brookings and BRAC, and the academic institutions included Brown, Duke, Georgetown, the London School of Economics, Princeton, Stanford, the University of Chicago, UC Berkeley, and UC San Diego. The survey provided a precise description of the intervention and asked respondents to predict the corresponding treatment impacts for several key outcomes we report here; see  A8 for the full text of the survey prompt.
Table 4 compares the actual values of the intervention’s components (column 1) to the respondents’ predicted values (column 2), and shows the $p$-value of a test of equality (column 3). Across all four components, differences between expert predictions and the values from our data are statistically significant. These respondents predicted, on average, that only 43% of the women in our study would be able to use digital payments (SD = 25.67pp). In practice, as described above, 99.75% of women purchased goods at least once. They also predicted, on average, that local authorities would attempt to tax roughly 40% of beneficiaries (SD = 27.28). In practice, less than 2% of the treatment group reported any diversion attempts, with no statistically significant difference in the control group. Moreover, respondents had little agreement in their views, reflected in the large variance across predictions. A combination of skepticism and uncertainty about potential efficacy might partly explain an unwillingness to adopt innovations, especially when stakes are high. Experts did a better job of predicting cost-efficiency, guessing on average that delivery costs would be 10.65 cents on the dollar, while the actual delivery costs were 6.7 cents on the dollar, indicating a more common belief that digital aid delivery is relatively cheap.
@@ 5 Discussion
We expect that whether the findings obtained from this context and population generalize to other settings will depend most critically on three key enabling conditions. First, vulnerable households must have access to phones. Second, there must be enough merchants who accept digital payments to allow convenient and competitive shopping. And third, markets must have enough goods to meet demand. Transitioning from in-kind to cash and digital support can increase competition for goods and so may increase market prices and thereby displace hunger to other vulnerable groups. Studies of famine highlight that in extreme cases of sudden and widespread deprivation – as in Gaza at the time of writing (Chotiner, 2024) – increasing food supply is critical, but that response to most famines historically required complementary strategies to redistribute resources (Dreze and Sen, 1990; Sen, 1999).
@@
Based largely on this research, humanitarian cash-based programming is growing and digital delivery is increasingly popular (Urquhart et al., 2023), but research during humanitarian crises with populations facing extreme cultural and political constraints is still nascent. In Online Appendix 1, we review the literature and find 23 RCTs of cash for humanitarian applications, 10 such RCTs in contexts categorized as “not free” by Freedom House, three of those 10 having some digital component, and none, that we can find, targeting populations that the government actively oppresses. This knowledge gap is consequential as these contexts and groups are increasingly affected by acute hunger (Foreign, Commonwealth and Development Office, 2023). As our expert survey demonstrates, our results are not readily inferred from prior studies: illiterate Afghan women could have been unable to utilize digital payments, local merchants could have refused to accept digital transactions, and the Taliban authorities could have interfered. Our study thus provides an “existence result” on the feasibility of digital aid in humanitarian response, with insight into the organizational and technological factors that contributed to success.
@@
The exigencies of this context and the fact that we work with a hard-to-reach group, however, necessitated three design choices that affect our generalizability. First, our experimental data cover only two months before the early group is treated. This time frame, however, is consistent with our learning objective: can digital payments cost-effectively reduce acute hunger. As such, we pre-specified this as the relevant time frame, and relatedly that we did not expect changes in outcomes like income or employment. Such impacts were unlikely – and did not realize – given the large humanitarian needs, prohibitions on women’s employment, and relatively modest transfer size. Second, working with our sample required approval of the de facto authorities. Our partners could not justify seeking Taliban permission to onboard women to the digital payments app without eventually providing humanitarian assistance. We thus opted for a short, staggered program where every participant received some assistance during the lean season. While the control group knew they would receive transfers, they were unaware when these would start and we find no evidence of borrowing or other anticipatory changes in behavior. Our late group, thus, provides a valid counterfactual. Third, because we could not visit our participants, and to avoid survey fatigue over the phone, we focused on a small set of questions and could not take measurements that would yield deeper insights (such as anthropometrics). Thus, other relevant topics (e.g., within household dynamics, validated mental health measures, and impacts on other household members) were not covered.
@@
Nonetheless, this exercise provides proof-of-concept that digital aid represents a potential cost-effective complement to existing modalities. We see five important directions for future research to help humanitarians select strategies to address hunger. First, relative to physical goods and cash, digital payments enable humanitarians to more easily vary when aid is delivered, its frequency, and the amount. Research can help identify how to optimally take advantage of this added flexibility. Second, more evidence is needed on the costs and benefits of digital relative to cash delivery. Some prior work addresses this question (Aker et al., 2016), but we think more of this work in fragile contexts controlled by hostile actors and where the vulnerable are hard-to-reach will be relevant given the rise in hunger among such groups. This is best pursued in experiments which randomize both digital delivery and cash delivery against a control group. Third, humanitarians need further evidence on how key enabling conditions – such as mobile phone penetration and the availability of merchants who accept digital payments – affect these relative costs and benefits, and how best to organize future hybrid partnerships based on differences in enabling conditions. Fourth, additional focus is needed on approaches to targeting, onboarding and monitoring beneficiaries in hostile environments that minimize risks to participants (Aiken et al., 2022; Jeong and Trako, 2022). In particular, the one point of physical contact in our study between humanitarians and beneficiaries was during onboarding. Further innovations might eliminate the need for contact between humanitarians and those in need altogether and apply automated screening of data from digital transactions and phone surveys to monitor user experience and identify potential fraudulent patterns. And fifth, implementing digital payments at scale may affect diversion strategies, market prices, and the broader use of digital payments technologies for other financial transactions. It may also spur the transformation of humanitarian organizations into leaner organizational models to best complement the capacities of private and community organizations. Future work with large-scale experiments should explore such changes.
Encouragingly, aid agencies are increasing their use of digital payments. The WFP is using insights from our study to scale digital payments to over 100,000 households in Afghanistan in partnership with HesabPay, which has been able to maintain its license to operate despite the dynamic strategic and political conditions in the country. As this program scales, it will need to grapple with the complex interdependencies between humanitarian agencies, digital payment platforms, and governments in fragile states. While local authorities could still choose to block digital aid provision, doing so might impede future humanitarian commitments and draw increased public attention to the issues of aid diversion and vulnerable populations. The centralized management of digital payments relative to cash or food distribution also reduces the need to engage with local actors who might complicate aid delivery and increases humanitarians’ bargaining power. The existence of key enablers for the success of a program like this in other fragile settings (e.g., the presence of CDCs, mobile phone availability, and digital payment platforms), as shown in A9, where Afghanistan ranks as a particularly challenging case, suggests that this approach could increase food security in other fragile states with high levels of hunger.
@@
(a) Skipped Meals
(b) Total Bread and Tea Meals
(c) Life Satisfaction
(d) Happy
Figure 2: Effects over Entire Duration of Program
@@
Control Control Treatment Standard Naive Adjusted Mean SD Effect Error p-value p-value N (1) (2) (3) (4) (5) (6) (7) Panel A. Primary Outcomes Days skipping meals (past week) 2.569 1.717 –0.76 0.051 0 0.0002 4761 Children skipping meals (=1) 0.873 0.333 –0.117 0.012 0 0.0002 4750 Regularly eat twice a day 0.501 0.5 0.093 0.015 0 0.0002 4736 Total bread and tea meals (past week) 13.639 3.549 –1.608 0.121 0 0.0002 4763 Food Security - KLK Index 0 1 0.501 0.032 0 4763 Better economic situation 0.048 0.213 0.335 0.011 0 0.0002 4762 Satisfied with fin. situation 0.133 0.34 0.263 0.012 0 0.0002 4755 Happy 0.154 0.361 0.28 0.014 0 0.0002 4732 Life satisfaction 3.179 1.707 1.963 0.068 0 0.0002 4763 Economic/Wellbeing - KLK Index 0 1 1.498 0.042 0 4763 Panel B. Secondary Outcomes Days eating rice (past week) 0.698 1.058 0.597 0.035 0 0.001 4763 Days eating beans (past week) 0.518 0.836 0.493 0.029 0 0.001 4763 Days eating vegetables (past week) 1.394 1.346 –0.003 0.041 0.934 0.453 4763 Days eating chicken (past week) 0.021 0.147 0.012 0.006 0.035 0.037 4763 Days eating dairy (past week) 0.074 0.383 0.047 0.013 0.001 0.002 4763 Able to buy medicine 0.051 0.221 0.034 0.01 0.001 0.002 3582 Involved in fin. decisions 0.666 0.472 0.017 0.015 0.272 0.185 4757 Total household income (past month) 876.683 1581.531 128.482 90.93 0.158 0.118 4763 Household’s head employed (past month) 0.192 0.394 –0.008 0.014 0.578 0.311 4741
@@
(1) (2) (3) (4) (5) Panel A. Food Security Days Skipping Children Skipping Regularly Eat Total Bread KLK Index Meals Meals (=1) Twice a Day (=1) & Tea Meals $\beta_{1}$: Treated $\times$ Round 2 –0.580*** –0.096*** –0.005 –0.948*** 0.309*** (0.094) (0.021) (0.028) (0.219) (0.056) $\beta_{2}$: Round 2 –0.195*** 0.025* 0.072*** –0.322** 0.096** (0.072) (0.013) (0.020) (0.151) (0.038) $\beta_{3}$: Treated –0.470*** –0.069*** 0.096*** –1.134*** 0.346*** (0.072) (0.015) (0.020) (0.162) (0.042) Control Mean 2.569 0.873 0.501 13.639 0.000 Observations 4,761 4,750 4,736 4,763 4,763 ($\beta_{1}$ + $\beta_{2}$ + $\beta_{3}$) / $\beta_{3}$ 2.650 2.040 1.700 2.120 2.170 Panel B. Wellbeing Better Economic Satisfied with Happy Life Satisfaction KLK Index Situation Financial Sit. $\beta_{1}$: Treated $\times$ Round 2 0.219*** 0.200*** 0.102*** 0.842*** 0.837*** (0.021) (0.024) (0.026) (0.127) (0.078) $\beta_{2}$: Round 2 –0.012 –0.036*** 0.030* –0.223*** –0.073* (0.009) (0.014) (0.017) (0.075) (0.040) $\beta_{3}$: Treated 0.226*** 0.163*** 0.229*** 1.542*** 1.079*** (0.015) (0.017) (0.019) (0.093) (0.055) Control Mean 0.048 0.133 0.154 3.179 0.000 Observations 4,762 4,755 4,732 4,763 4,763 ($\beta_{1}$ + $\beta_{2}$ + $\beta_{3}$) / $\beta_{3}$ 1.910 2.010 1.580 1.400 1.710
$Notes$: This table reports estimated impacts of treatment separately for the first and second survey round. Households were surveyed once per month for two months. Each of these months constitutes a survey round. All specifications control for stratum fixed effects and the baseline value of the dependent variable, if available. Standard errors are clustered at individual level. The outcome variables follow the primary outcomes shown in Table 1.
@@
Gov. Off. Comm. Leader Gov. Off. Comm. Leader KLK Yes to Others Others You You Index Any Question (1) (2) (3) (4) (5) (6) Panel A. Baseline Treated 0.001 0.002 0.002 0.002 0.074* 0.003 (0.001) (0.002) (0.001) (0.003) (0.042) (0.003) Observations 4,558 4,596 4,631 4,626 4,648 4,509 Control Mean 0.002 0.002 0.001 0.007 0.012 0.009 Panel B. Long-Run $\beta_{1}$: Treated $\times$ Round 2 –0.000 0.006* 0.002 –0.001 0.091 0.004 (0.003) (0.003) (0.003) (0.005) (0.084) (0.006) $\beta_{2}$: Round 2 –0.002 –0.003 –0.003* –0.008** –0.137*** –0.010** (0.002) (0.002) (0.002) (0.003) (0.045) (0.004) $\beta_{3}$: Treated 0.001 –0.007 –0.001 0.003 –0.062 –0.004 (0.005) (0.005) (0.005) (0.009) (0.140) (0.011) ($\beta_{1}$ + $\beta_{2}$ + $\beta_{3}$) / $\beta_{3}$ –0.770 0.500 1.880 –1.680 1.740 2.550 Observations 4,558 4,596 4,631 4,626 4,648 4,509 Panel C. List Experiment All Sample Late Sample Early Sample Treated List 0.011 0.019 0.008 (0.023) (0.032) (0.032) Observations 2,358 1,171 1,187 Mean Items Mentioned 0.746 0.691 0.800
@@
Table 4: Actual Values vs. Experts’ Predictions
Variable Actual Values Predicted Values p-value Predicted = Actual (1) (2) (3) Percent Able to Use Digital Payments 99.75 43.82 $<$ 0.001 (2.75) Percent Reporting Diversion Attempts 1.99 39.85 $<$ 0.001 (2.98) Delivery Costs (cents per \$1 delivered) 6.7 10.65 $<$ 0.001 (0.74) How Many Bread and Tea Meals in Past Week? 11.96 10.32 $<$ 0.001 (0.45)